40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 30 00:02:25 2013, Jenne, Update, LSC, LSC demod boards need some thought 
    Reply  Wed Oct 30 01:22:56 2013, Jenne, Update, LSC, REFL 165 demod phase adjusted 
       Reply  Wed Oct 30 01:44:13 2013, Jenne, Update, LSC, MICH and PRCL gains adjusted (Config file saved) 
          Reply  Wed Oct 30 03:36:51 2013, Jenne, Update, LSC, MICH and PRCL UGFs change with ALS enabled 
             Reply  Mon Nov 4 15:46:17 2013, Jenne, Update, LSC, Thoughts and Conclusions from last week's PRMI+2arms attempt 12x
                Reply  Tue Nov 5 19:39:27 2013, Jenne, Update, LSC, OpLev time series OpLevs_during_PRMI_2arms.pdf
    Reply  Wed Oct 30 03:33:17 2013, Rana, Update, LSC, LSC demod boards need some thought 0309.png
       Reply  Wed Oct 30 08:32:55 2013, Evan, HowTo, Computer Scripts / Programs, matplotlibrc 
Message ID: 9349     Entry time: Tue Nov 5 19:39:27 2013     In reply to: 9338
Author: Jenne 
Type: Update 
Category: LSC 
Subject: OpLev time series 

[Rana, Jenne]

We looked at the time series for all the oplevs except the BS, from last Tuesday night, during a time when we were building up the power in the arms.  We conclude from a 400 second stretch of data that there is not discernible difference in the amount of motion of any optic, when the cavities are at medium power, and when they're at low power.  Note however, that we don't have such a nice stretch of data for the really high powers, so the maximum arm power in these plots is around 5.  Both the TRX and TRY signals look fairly stationary up to powers of 1 or 2, but once you get to 4 or 5, the power fluctuations are much more significant.  So, since this isn't caused by any optic moving more, perhaps it's just that we're more sensitive to optic motion when we're closer to resonance in the arms.

However, from this plot, it looks like the ETMY is moving much more than any other optic.  On the other hand, ETMY has not ever been calibrated (there's an arbitrary 300 in there for the calibration numbers on the ETMY oplev screen).  So, perhaps it's not actually moving any more than other optics.  We should calibrate the ETM oplevs nicely, so we have some real numbers in there.  ETMX also only is roughly calibrated, relative to the OSEMs.  We should either do the move-the-QPD calibration, or a Kakeru-style pitch and yaw some mirrors and look at transmitted power.

Traces on this xml file have been filtered with DTT, using zpk([0],[0.03],1,"n").

OpLevs_during_PRMI_2arms.pdf

 

ELOG V3.1.3-